OpenStudy (anonymous):

Is the simplified form of 2square root of 3 • square root of 12 rational?

OpenStudy (haseeb96):

2sqrt3 multiply with sqrt 12 is it yur question?

OpenStudy (anonymous):

First, multiply it out. Then, what is in the square root?

OpenStudy (anonymous):

\[2\sqrt{3}*\sqrt{12}\]

OpenStudy (anonymous):

i then got \[2\sqrt{3}*2\sqrt{3}\]

OpenStudy (haseeb96):